OpenStudy (anonymous):

Need help guys plss solve this and show your solutions : Problem : Find the derivatives of the function . y=(2x-7)(x+3)

OpenStudy (anonymous):

multiply it out 2x^2-x-21 take derivative 2*2x-1 = 4x-1

OpenStudy (anonymous):

^ that or you can use the product rule which is y' = f'g+ fg', if they are tell you to do it with that specific method

OpenStudy (anonymous):

the answer is wrong . need more help plss. Our topic is definition of derivatives

zepdrix (zepdrix):

So you need to use the limit definition for this problem? \[\huge f'(x)=\lim_{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Oh boy :\

zepdrix (zepdrix):

Hmm well I still think multiplying them out first is going to be the easier way, but not by much hehe.

zepdrix (zepdrix):

Write y as a function of x, it might make this a little easier on the notation. \[\huge f(x)=2x^2-x-21\]\[\huge f(x+h)=2(x+h)^2-(x+h)-21\]

zepdrix (zepdrix):

And then plugging it in, you get something like this... and need to do a bunch of simplification :3 oh boy!\[\large \lim_{h \rightarrow 0} \frac{2(x+h)^2-(x+h)-21-(2x^2-x-21)}{h}\]

OpenStudy (anonymous):

answer : 4x ??

zepdrix (zepdrix):

Hmm there shoulda been a 1 in there somewhere :D I think you'll get 4x-1 somehow.